OpenStudy (anonymous):

Determine two coterminal angles in radian measure (one positive and one negative) for each angle. a. 7pi/8 b. 8pi/45

OpenStudy (luigi0210):

Do you know what a co-terminal angle is?

OpenStudy (anonymous):

not at all :O

OpenStudy (luigi0210):

Maybe we should figure that out first then xD

OpenStudy (anonymous):

that sounds like a plan! :p

OpenStudy (luigi0210):

But it's basically adding/subtracting \(2 \pi/360^○\) http://hotmath.com/hotmath_help/topics/coterminal-angles.html

OpenStudy (anonymous):

so would it be 7pi +2pi first?

OpenStudy (luigi0210):

\[\LARGE \frac{7 \pi}{8}\pm2 \pi\] Yes~

OpenStudy (anonymous):

I got 23pi/8 and -9pi/8... sound about right?

OpenStudy (luigi0210):

And you would be correct Miss Ali~

OpenStudy (luigi0210):

Now do the same for the second.

OpenStudy (anonymous):

so that would be for a. so b. would be -82pi/45 and 98pi/45?

OpenStudy (luigi0210):

And you're done :)

OpenStudy (anonymous):

thanks so much!!! :D